find the slope if it exsists of the line containing the pair of points
(-3.3,15.5) and (-0.8,5.6)

Found 2 solutions by subudear, rfer:
Answer by subudear(62)

slope of line is given by m = (y2-y1)/(x2-x1)
given points are
(-3.3,15.5) and (-0.8,5.6)
x1 = -3.3, y1 = 15.5
x2 = -0.8, y2 = 5.6
replacing the values we get
m = (5.6 - 15.5)/(-0.8 - (-3.3))
m = -3.96

Answer by rfer(16322)

slope m= (5.6-15.5)/(-0.8--3.3)
m=-9.9/2.5
m=-4 round off from -3.96